# Personalized Attacks of Social Engineering in Multi-turn Conversations - LLM Agents for Simulation and Detection

### v1

@@ Abstract
The rapid advancement of conversational agents, particularly, chatbots powered by Large Language Models (LLMs), poses a significant risk of social engineering (SE) attacks on social media platforms. SE detection in multi-turn, chat-based interactions is considerably more complex than single-instance detection due to the dynamic nature of these conversations. A critical factor in mitigating this threat is understanding the mechanisms through which SE attacks operate, specifically how attackers exploit vulnerabilities and how victims’ personality traits contribute to their susceptibility. In this work, we propose an LLM-agentic framework, SE-VSim, to simulate SE attack mechanisms by generating multi-turn conversations. We model victim agents with varying personality traits to assess how psychological profiles influence susceptibility to manipulation. Using a dataset of over 1,000 simulated conversations, we examine attack scenarios in which adversaries, posing as recruiters, funding agencies, and journalists, attempt to extract sensitive information. Based on this analysis, we present a proof of concept, SE-OmniGuard to offer personalized protection to users by leveraging prior knowledge of the victim’s personality, evaluating attack strategies, and monitoring information exchanges in conversations to identify potential SE attempts. Our code and data are available at following repository.
Personalized Attacks of Social Engineering in Multi-turn Conversations -
LLM Agents for Simulation and Detection
Tharindu Kumarage1, Cameron Johnson2, 11footnotemark: 1, Jadie Adams2,thanks: These authors contribute to this work equally., Lin Ai3, 11footnotemark: 1, Matthias Kirchner2, Anthony Hoogs2, Joshua Garland1, Julia Hirschberg3, Arslan Basharat2, Huan Liu1 1Arizona State University, 2Kitware, Inc, 3Columbia University, {kskumara, joshua.garland, huanliu}@asu.edu {jadie.adams, cameron.johnson, arslan.basharat}@kitware.com
@@ 1 Introduction
The growing sophistication of conversational agents, especially those powered by Large Language Models (LLMs), presents a major risk for misuse in social engineering (SE) attacks across digital communication platforms Schmitt and Flechais (2023). LLM-powered SE represents a significant threat, as these models can produce highly convincing, human-like interactions in real-time, greatly increasing the success rate of attacks. Unlike traditional SE, which often reveals itself through signs like grammatical errors or implausible scenarios, LLM-based attacks generate coherent, contextually relevant dialogues that are more difficult to detect. Detecting SE in multi-turn, chat-based interactions is especially challenging due to the dynamic nature of these conversations, where the interaction evolves with each exchange.
@@
In recent years, the application of LLMs as world simulators has gained traction, with numerous studies utilizing LLMs to emulate sophisticated cyber-attacks in an effort to develop effective defense mechanisms against potential future threats Xu et al. (2024); Wang et al. (2024). A recent study conducted such a simulation of LLM-powered SE attacks, discussing the dual role of LLMs as both a perpetrator and a defender in chat-based-SE (CSE) scenarios Ai et al. (2024). While these duality-based simulations represent an important first step towards protecting users from LLM-powered CSE attacks, further considerations are necessary to effectively ground these simulations and defense mechanisms in real-world CSE contexts.
@@
We identify two key limitations in both the LLM simulation of CSE and the LLM’s defense mechanisms: (1) the lack of grounding in conceptual frameworks for SE attack mechanisms Wang et al. (2021)—specifically, how attackers exploit vulnerabilities and how victims’ personality traits contribute to these susceptibilities. Without this grounding, the simulated conversations may diverge significantly from real-world scenarios; and (2) the overemphasis on detecting sensitive information exchange as the primary indicator of a successful CSE attack. In reality, successful CSE attacks may not immediately involve the exchange of sensitive information; instead, attacks often begin by building trust with the victims, laying the groundwork for more severe attacks in the future Salahdine and Kaabouch (2019).
@@
To address these limitations, we propose an agentic simulation, SE-VSim, designed to emulate realistic CSE attacks by independently modeling both an attacker and victim agent, grounded in concepts from SE effect mechanisms Wang et al. (2021). The victim agent is modeled with varying psychological profiles based on the Big Five personality traits (openness, conscientiousness, extraversion, agreeableness, and neuroticism) Goldberg (2013); Cusack and Adedokun (2019), enabling exploration of how different personality traits influence vulnerability to SE attacks. Using this framework, we generate a high-quality dataset of 1,350 simulated conversations that represent real-world CSE scenarios, where the attacker poses as a recruiter, funding agency, or journalist, attempting to extract sensitive information such as personally identifiable information (PII), financial data, or intellectual property.
@@
Based on these conversations, we argue that an ideal defense should not only focus on identifying sensitive information exchanges but consider the nuances introduced by victim personality traits and attack strategies. To this end, we demonstrate a proof of concept, SE-OmniGuard, which incorporates prior knowledge of the victim’s personality, evaluates attacker strategies, and monitors information exchanges throughout the conversation to identify potential SE attempts, thereby offering personalized protection to users.
@@
  • A proof of concept, SE-OmniGuard: Vision for a defense that incorporates victim personality traits, monitors attack strategies and evaluates conversation dynamics to detect SE attempts, providing personalized protection.
@@ 2 Preliminaries
The Big Five Personality Traits (Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism) serve as a robust, cross-culturally validated framework for modeling psychological profiles in a consistent mannerGoldberg (2013); Cusack and Adedokun (2019). Studies in cybersecurity and phishing research have demonstrated that these traits significantly influence an individual’s susceptibility to SE attacks Rahman et al. (2022). For example, High agreeableness is associated with an increased tendency to comply with requests, low conscientiousness often aligns with less guarded online behaviors, and high neuroticism can manifest as a heightened willingness to engage due to anxiety or urgency—factors that attackers frequently exploit López-Aguilar and Solanas (2021); Bright et al. (2022). Beyond their empirical validation in SE contexts, the Big Five model stands out due to its broad applicability across diverse populations and psychological constructs Costa Jr and McCrae (1992); McCrae and Terracciano (2005). It offers a structured means to describe the victim personas through natural language descriptions, making it straightforward to incorporate them into LLM-based simulations.
@@ 3 Related Work
SE attacks commonly occur through communication channels such as SMS, phone calls, and online platforms, including social media Tsinganos et al. (2018); Zheng et al. (2019). Researchers have studied the phases of SE attacks extensively Zheng et al. (2019); Wang et al. (2021); Karadsheh et al. (2022), leading to various defense mechanisms. For example, SEADER++ Lansley et al. (2020) detects malicious chats using synthetic datasets and an MLP classifier, while ICSA Yoo and Cho (2022) employs TextCNN-based classifiers to address phishing stages on social networks. Recent advancements include fine-tuned models like SG-CSE BERT Tsinganos et al. (2023) for zero-shot SE detection and CSE-ARS Tsinganos et al. (2024), a late-fusion approach combining multiple models to enhance detection across contexts.
@@
Individual personality traits pose challenges for designing effective phishing defenses Anawar et al. (2019), as they influence susceptibility to manipulation Rahman et al. (2022). Studies using models like the Big Five Cusack and Adedokun (2019) show that traits such as high agreeableness or low conscientiousness increase vulnerability to phishing and deception Cho et al. (2016); Anawar et al. (2019). While integrating personality recognition into defense systems improves detection An et al. (2018), most SE defenses still adopt a one-size-fits-all approach. To address this, our work introduces personality-aware simulations using LLMs Hu and Collier (2024); Schuller et al. (2024); Sun et al. (2024) to explore how psychological traits influence susceptibility to manipulation.
LLM Agents and Cyber-Attacks
Traditional SE defenses focus on human-initiated attacks, but the rise of LLMs introduces new threats. LLMs mimic human conversational patterns, trust cues Mireshghallah et al. (2024); Hua et al. (2024), and elicit emotions Miyakawa et al. (2024); Gong et al. (2023), enabling sophisticated digital deception Wu et al. (2024); Ai et al. (2021, 2023); Kumarage et al. (2024); Beigi et al. (2024) and SE attacks Schmitt and Flechais (2023). While efforts exist to simulate cyberattacks using LLMs Xu et al. (2024); Happe and Cito (2023); Kumarage et al. (2023); Naito et al. (2023); Fang et al. (2024), LLM-driven SE attacks remain underexplored. Asfour and Murillo (2023) modeled human responses to SE attacks via LLMs, but comprehensive multi-turn conversational frameworks are lacking. Ai et al. (2024) advanced the field by exploring LLMs as both enablers and defenders against SE attacks but overlooked the role of victim personality traits. Our work addresses this gap by integrating personality-aware defense strategies for dynamic, personalized protection against LLM-powered SE threats.
@@ 4 Simulating Social Engineering Effect Mechanisms
This section outlines our framework, SE-VSim, designed to simulate SE effect mechanisms. The goal is to model multi-turn conversations between an attacker and a victim agent grounded in a real-SE conceptual framework Wang et al. (2021). By modeling the interaction between attack strategies and victim vulnerabilities, we aim to explore how personality traits influence susceptibility to SE attacks. As shown in Figure 1, the framework consists of three key components: the attacker agent, the victim agent, and a conversation generation pipeline that enables dynamic interactions between these agents. Both the attacker and victim agents are implemented using open-source LLMs.
Figure 1: Components of the framework, SE-VSim
@@ 4.1 Attacker Agent
The attacker agent is designed to emulate a malicious actor in a multi-turn SE scenario. To simulate this behavior, we condition the attacker agent’s intent through in-context learning using a predefined attack goal $G_{se}$. The $G_{se}$ consists of two parts: (i) role $A_{role}$ the role the attacker is pretending to be, and (ii) attack intent, $A_{intent}$ defines the malicious goal, i.e., extract a piece of target information from the victim. As mentioned in the Preliminaries, we focus on attacker roles that commonly appear in LinkedIn-type professional contexts: Funding Agencies (AF), Journalists (JO), and Recruiters (RE). Corresponding to these roles, we define three target information types: Personally Identifiable Information (PII), sensitive financial details, and intellectual property (patents and trademarks).
@@
Where $C_{\text{prior}}^{i-1}$ represents the context of the previous conversation turns and $G_{\text{se}}=A_{intent}\oplus A_{role}$ represents the malicious SE goal, such as extracting sensitive financial information pretending to be a journalist. The prompts used for the attacker agent can be found in the appendix A. In order to generate benign conversations, we remove the malicious attack intent from the agent’s goal, $G_{\text{benign}}=G_{\text{se}}-A_{intent}$.
@@ 4.2 Victim Agent
The victim agent is designed to represent individuals with varying personality traits, affecting their vulnerability to SE attacks. We model the victim’s psychological profile based on the Big Five personality traits. Detailed natural language descriptions of each personality trait can be found in the appendix Table 6. Each victim agent’s persona is conditioned through in-context learning to exhibit specific personality-driven responses during conversations. Formally, the victim’s response at each turn $i$ can be modeled as:
@@
Where $C_{\text{prior}}^{i-1}$ represents the context of the previous conversation turns and $P_{\text{trait}}$ is the context representing the victim’s personality traits. This conditioning allows us to explore how different personality traits influence the victim’s susceptibility to manipulation, which consequently increases the diversity of the simulated conversations. For instance, a highly agreeable victim might be more trusting, while a more neurotic individual might respond with suspicion, influencing the attacker’s approach and the eventual outcome of the SE attempt. The prompts used for the victim agent can be found in the appendix Table 4.
@@ 4.3 Conversation Generation
Where $C_{\text{prior}}^{t+1}$ is the updated conversation state at turn $t+1$. Implementation details of the overall framework, including LLM generation parameters, can be found in the appendix A.
@@
Here $L_{\text{intent}}$ represents the overall maliciousness label for the conversation. $G_{\text{se}}$ is the attacker’s goal involving the extraction of sensitive information or manipulation, resulting in a malicious label. $G_{\text{benign}}$ represents a neutral goal, leading to a benign conversation label.
@@ 4.4 Conversation Statistics
Incorporating the SE-VSim framework, we generate a dataset consisting of 1,350 conversations, categorized into malicious (900 conversations with malicious intent in the attacker agent instructions) and benign (450 conversations with neutral intent in the instructions) interactions. The conversations are further divided based on the attacker roles: Funding Agencies (AF), Journalists (JO), and Recruiters (RE). Each scenario contains 100 malicious and 50 benign conversations per target information type, which includes Personally Identifiable Information (PII), financial sensitive information, and patents and trademarks. Moreover, Within each subset of conversations, 10 conversations per victim traits are represented, ensuring a diverse set of victim profiles. Figure 2 illustrates the distribution of conversations, demonstrating the number of conversations across attacker roles, target information types, and victim traits.
Figure 2: Dataset Distribution. Left: Proportion of malicious and benign conversations. Right: Proportion information types and attacker roles within both malicious and benign conversations.
@@ 4.5 Conversation Annotation
In addition to the general conversation labels (malicious or benign), in our work, we provide two additional annotations. Primarily, we evaluate the successfulness of the malicious conversations using a 3-level metric. This metric measures how well the attacker achieves their goal, whether by obtaining sensitive information or by gaining the victim’s trust for future manipulation. We denote the success of the conversation $S_{\text{success}}$ as:
@@
To this end, we introduce GPT-4o-mini as a third annotator, assessing whether it could match human performance in classifying attack success. We computed Fleiss’ Kappa agreement ($k=0.796$), indicating a high level of alignment between human annotators and the LLM. These results suggest that, in the context of CSE annotation, a well-instructed LLM can provide reliable success labels, reducing the burden on human annotators without compromising accuracy. For further details on the annotation guidelines, including the criteria for assigning success scores, please refer to the appendix B.
A secondary annotation task focuses on labeling attack strategies—the tactics used by attackers to manipulate victims. Inspired by established social engineering frameworks Wang et al. (2021), we categorized these strategies into high-level tactics, including: Persuasion, Social Influence, and Cognition, Attitude, and Behavior. Each high-level category consists of sub-categories that describe specific manipulative techniques used in social engineering attempts. Unlike attack success annotation, attack strategy annotation is supplementary and involves complex, multi-label classification across thousands of messages. Given budget constraints, manual annotation of such granular labels is not feasible. Instead, we rely on GPT-4o-mini as an LLM judge to identify and categorize attack strategies. For further details regarding the attack strategy definitions, and annotation guidelines, please refer to Appendix B.

### v2

# Personalized Attacks of Social Engineering in Multi-turn Conversations: LLM Agents for Simulation and Detection

@@ Abstract
The rapid advancement of conversational agents, particularly chatbots powered by large language models (LLMs), poses a significant risk of social engineering (SE) attacks on social media platforms. SE detection in multi-turn, chat-based interactions is considerably more complex than single-instance detection due to the dynamic nature of these conversations. A critical factor in mitigating this threat is understanding SE attack mechanisms, specifically how attackers exploit vulnerabilities and how victims’ personality traits contribute to their susceptibility. In this work, we propose an LLM-agentic framework, SE-VSim, to simulate SE attack mechanisms by generating multi-turn conversations. We model victim agents with varying personality traits to assess how psychological profiles influence susceptibility to manipulation. Using a dataset of over 1,000 simulated conversations, we examine attack scenarios in which adversaries posing as recruiters, funding agencies, and journalists attempt to extract sensitive information. Based on this analysis, we present a proof of concept, SE-OmniGuard, to offer personalized protection to users by leveraging prior knowledge of the victim’s personality, evaluating attack strategies, and monitoring information exchanges in conversations to identify SE attempts. Our code and data are available at following repository.
footnotetext: Distribution Statement: Approved for Public Release, Distribution Unlimited.
@@ 1 Introduction
The growing sophistication of conversational agents, especially those powered by large language models (LLMs), presents a major risk for misuse in social engineering (SE) attacks across digital communication platforms Schmitt & Flechais (2023). LLM-powered SE represents a significant threat, as these models can produce highly convincing, human-like interactions in real time, greatly increasing the success rate of attacks. Unlike traditional SE, which often reveals itself through signs like grammatical errors or implausible scenarios, LLM-based attacks generate coherent, contextually relevant dialogues that are more difficult to detect. Detecting SE in multi-turn, chat-based interactions is especially challenging due to the dynamic nature of these conversations, where the interaction evolves with each exchange.
@@
In recent years, the application of LLMs as world simulators has gained traction, with numerous studies utilizing LLMs to emulate sophisticated cyberattacks to develop effective defense mechanisms against future threats Xu et al. (2024); Wang et al. (2024). A recent study conducted such a simulation of LLM-powered SE attacks, discussing the dual role of LLMs as both a perpetrator and a defender in chat-based SE (CSE) scenarios Ai et al. (2024). While these duality-based simulations represent an important first step towards protecting users from LLM-powered CSE attacks, further considerations are necessary to effectively ground these simulations and defense mechanisms in real-world CSE contexts.
@@
We identify two key limitations in both the LLM simulation of CSE and the LLM’s defense mechanisms: (1) the lack of grounding in conceptual frameworks for SE attack mechanisms Wang et al. (2021)—specifically, how attackers exploit vulnerabilities and how victims’ personality traits contribute to these susceptibilities. Without this grounding, the simulated conversations may diverge significantly from real-world scenarios, and (2) the overemphasis on detecting sensitive information exchange as the primary indicator of a successful CSE attack. In reality, successful CSE attacks may not immediately involve the exchange of sensitive information; instead, attacks often begin by building trust with the victims, laying the groundwork for more severe attacks in the future Salahdine & Kaabouch (2019).
@@
To address these limitations, we propose an agentic simulation, SE-VSim, designed to emulate realistic CSE attacks by independently modeling both an attacker and victim agent, grounded in concepts from SE effect mechanisms Wang et al. (2021). The victim agent is modeled with varying psychological profiles based on the Big Five personality traits (openness, conscientiousness, extraversion, agreeableness, and neuroticism) Goldberg (2013); Cusack & Adedokun (2019), enabling exploration of how different personality traits influence vulnerability to SE attacks. Using this framework, we generate a high-quality dataset of 1,350 simulated conversations that represent real-world CSE scenarios, where the attacker poses as a recruiter, funding agency, or journalist, attempting to extract sensitive information.
@@
Based on these conversations, we argue that an ideal defense should not only focus on identifying sensitive information exchanges but also consider the nuances introduced by victim personality traits and attack strategies. To this end, we demonstrate a proof of concept, SE-OmniGuard, which incorporates prior knowledge of the victim’s personality, evaluates attacker strategies, and monitors information exchanges throughout the conversation to identify potential SE attempts, thereby offering personalized protection to users.
@@
  • A proof of concept, SE-OmniGuard: Vision for a defense that incorporates victim personality traits, monitors attack strategies, and evaluates conversation dynamics to detect SE attempts, providing personalized protection. We will release the code and dataset upon acceptance or publication decision.
@@ 2 Preliminaries
The Big Five personality traits (Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism) serve as a robust, cross-culturally validated framework for modeling psychological profiles in a consistent mannerGoldberg (2013); Cusack & Adedokun (2019). Studies in cybersecurity and phishing research have demonstrated that these traits significantly influence an individual’s susceptibility to SE attacks Rahman et al. (2022). For example, high agreeableness is associated with an increased tendency to comply with requests, low conscientiousness often aligns with less guarded online behaviors, and high neuroticism can manifest as a heightened willingness to engage due to anxiety or urgency—factors that attackers frequently exploit López-Aguilar & Solanas (2021); Bright et al. (2022). Beyond their empirical validation in SE contexts, the Big Five model stands out due to its broad applicability across diverse populations and psychological constructs Costa Jr & McCrae (1992); McCrae & Terracciano (2005). It offers a structured means to describe the victim personas through natural language, making it straightforward to incorporate them into LLM-based simulations.
@@ 3 Related Work
SE attacks commonly occur through communication channels such as SMS, phone calls, and online platforms, including social media Tsinganos et al. (2018); Zheng et al. (2019). Researchers have studied the phases of SE attacks extensively Zheng et al. (2019); Wang et al. (2021); Karadsheh et al. (2022), leading to various defense mechanisms. For example, SEADER++ Lansley et al. (2020) detects malicious chats using synthetic datasets and an MLP classifier, while ICSA Yoo & Cho (2022) employs TextCNN-based classifiers to address phishing stages on social networks. Recent advancements include fine-tuned models like SG-CSE BERT Tsinganos et al. (2023) for zero-shot SE detection and CSE-ARS Tsinganos et al. (2024), a late-fusion approach combining multiple models to enhance detection across contexts.
@@
Individual personality traits pose challenges for designing effective phishing defenses Anawar et al. (2019), as they influence susceptibility to manipulation Rahman et al. (2022). Studies using models like the Big Five Cusack & Adedokun (2019) show that traits such as high agreeableness or low conscientiousness increase vulnerability to phishing and deception Cho et al. (2016); Anawar et al. (2019). While integrating personality recognition into defense systems improves detection An et al. (2018), most SE defenses still adopt a one-size-fits-all approach. To address this, our work introduces personality-aware simulations using LLMs Hu & Collier (2024); Schuller et al. (2024); Sun et al. (2024) to explore how psychological traits influence susceptibility to manipulation.
LLM Agents and Cyberattacks
Traditional SE defenses focus on human-initiated attacks, but the rise of LLMs introduces new threats. LLMs mimic human conversational patterns, trust cues Mireshghallah et al. (2024); Hua et al. (2024), and elicit emotions Miyakawa et al. (2024); Gong et al. (2023), enabling sophisticated digital deception Wu et al. (2024); Ai et al. (2021; 2023); Kumarage et al. (2024); Beigi et al. (2024) and SE attacks Schmitt & Flechais (2023). While efforts exist to simulate cyberattacks using LLMs Xu et al. (2024); Happe & Cito (2023); Kumarage et al. (2023); Naito et al. (2023); Fang et al. (2024), LLM-driven SE attacks remain underexplored. Asfour & Murillo (2023) modeled human responses to SE attacks via LLMs, but comprehensive multi-turn conversational frameworks are lacking. Ai et al. (2024) advanced the field by exploring LLMs as both enablers and defenders against SE attacks, but overlooked the role of victim personality traits. Our work addresses this gap by integrating personality-aware defense strategies for dynamic, personalized protection against LLM-powered SE threats.
@@ 4 Simulating Social Engineering Effect Mechanisms
This section outlines our framework, SE-VSim, designed to simulate SE effect mechanisms. The goal is to model multi-turn conversations between an attacker and a victim agent based on a realistic SE conceptual framework Wang et al. (2021). By modeling the interaction between attack strategies and victim vulnerabilities, we aim to explore how personality traits influence susceptibility to SE attacks. As shown in Figure 1, the framework consists of three key components: the attacker agent, the victim agent, and a conversation generation pipeline that enables dynamic interactions between these agents. Both the attacker and victim agents are implemented using open-source LLMs.
Figure 1: Components of the SE-VSim framework.
@@ 4.1 Attacker Agent
The attacker agent is designed to emulate a malicious actor in a multi-turn SE scenario. To simulate this behavior, we condition the attacker agent’s intent through in-context learning using a predefined attack goal $G_{se}$. The $G_{se}$ consists of two parts: (i) role $A_{role}$ - the role the attacker is pretending to be, and (ii) attack intent $A_{intent}$ - defines the malicious goal, i.e., extract a piece of target information from the victim. As mentioned in the Preliminaries, we focus on attacker roles that commonly appear in LinkedIn-type professional contexts: Funding Agencies (AF), Journalists (JO), and Recruiters (RE). Corresponding to these roles, we define three target information types: Personally Identifiable Information (PII), sensitive financial details, and intellectual property (patents and trademarks).
@@
Where $C_{\text{prior}}^{i-1}$ represents the context of the previous conversation turns and $G_{\text{se}}=A_{intent}\oplus A_{role}$ represents the malicious SE goal, such as extracting sensitive financial information by impersonating a journalist. The prompts used for the attacker agent are listed in Appendix B. To generate benign conversations, we remove the malicious attack intent from the agent’s goal, $G_{\text{benign}}=G_{\text{se}}-A_{intent}$.
@@ 4.2 Victim Agent
The victim agent is designed to represent individuals with varying personality traits, affecting their vulnerability to SE attacks. We model the victim’s psychological profile based on the Big Five personality traits. Detailed natural language descriptions of each personality trait can be found in Table 6 of the Appendix. Each victim agent’s persona is conditioned through in-context learning to exhibit specific personality-driven responses during conversations. Formally, the victim’s response at each turn $i$ can be modeled as:
@@
Where $C_{\text{prior}}^{i-1}$ represents the context of the previous conversation turns, and $P_{\text{trait}}$ is the context representing the victim’s personality traits. This conditioning allows us to explore how different personality traits influence the victim’s susceptibility to manipulation, which consequently increases the diversity of the simulated conversations. For instance, a highly agreeable victim might be more trusting, while a more neurotic individual might respond with suspicion, influencing the attacker’s approach and the eventual outcome of the SE attempt. The prompts used for the victim agent are listed in Table 4 of the Appendix.
@@ 4.3 Conversation Generation
Where $C_{\text{prior}}^{t+1}$ is the updated conversation state at turn $t+1$. Implementation details of the overall framework, including LLM generation parameters, can be found in Appendix B.
@@
Here $L_{\text{intent}}$ represents the overall maliciousness label for the conversation. $G_{\text{se}}$ is the attacker’s goal to extract sensitive information or manipulate, resulting in a malicious label. $G_{\text{benign}}$ represents a neutral goal, leading to a benign conversation label.
@@ 4.4 Conversation Statistics
Using the SE-VSim framework, we generate a dataset consisting of 1,350 conversations. The dataset includes 900 malicious conversations (with malicious intent in the attacker agent instructions) and 450 benign conversations (with neutral intent in the instructions). The conversations are further divided based on the attacker roles: Funding Agencies (AF), Journalists (JO), and Recruiters (RE). Each scenario contains 100 malicious and 50 benign conversations per target information type, which includes PII, sensitive financial information, and patents and trademarks. Moreover, within each subset of conversations, 10 conversations per victim trait are represented, ensuring a diverse set of victim profiles. Figure 2(a) illustrates the distribution of conversations, demonstrating the number of conversations across attacker roles, target information types, and victim traits.
(a) Dataset distribution: (Left) malicious vs. benign, (Right) info types and attacker roles.
(b) Attack success distribution over personality traits.
Figure 2: (a) Dataset characteristics and (b) personality-trait-wise attack success trends.
@@ 4.5 Conversation Annotation
In addition to the general conversation labels (malicious or benign), we provide two additional annotations in our work. Primarily, we evaluate the successfulness of the malicious conversations using a 3-level metric. This metric measures how well the attacker achieves their goal, whether by obtaining sensitive information or by gaining the victim’s trust for future manipulation. We denote the success of the conversation $S_{\text{success}}$ as:
@@
To this end, we introduce GPT-4o-mini as a third annotator, assessing whether it could match human performance in classifying attack success. We computed Fleiss’ Kappa agreement ($k=0.796$), indicating a high level of alignment between human annotators and the LLM. These results suggest that, in the context of CSE annotation, a well-instructed LLM can provide reliable success labels, reducing the burden on human annotators without compromising accuracy. For further details on the annotation guidelines, including the criteria for assigning success scores, refer to Appendix C.
A secondary annotation task focuses on labeling attack strategies—the tactics used by attackers to manipulate victims. Inspired by established social engineering frameworks Wang et al. (2021), we categorized these strategies into high-level tactics: Persuasion; Social Influence; and Cognition, Attitude, and Behavior. Each high-level category consists of sub-categories that describe specific manipulative techniques used in social engineering attempts. Unlike attack success annotation, attack strategy annotation is supplementary and involves complex, multi-label classification across thousands of messages. Given budget constraints, manual annotation of such granular labels is not feasible. Instead, we rely on GPT-4o-mini as an LLM judge to identify and categorize attack strategies. For further details regarding the attack strategy definitions and annotation guidelines, refer to Appendix C.